1+1′ TWO-COLOR PHOTOIONIZATION SPECTROSCOPY OF ACETYLENE

SHUENN-JIUN TANG, JIM JR-MIN LIN, KENNY NI, and \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P.O. Box 23-166, Taipei, Taiwan, R.O.C. and Department of Chemistry, National Taiwan University, Taipei, Taiwan, R.O.C..

1+1′ two-color photoionization spectrum of C_{2}H_{2} via a single rovibrational level of Ã^{1}A_{u} as the intermediate was recorded under nearly field-free condition in the range of 91900-92700 cm^{−1}. If the first color was frequency-tuned to excite a single rotational level (J′ = 1 − 4, e parity) of V_{4}K_{1} of the Ã^{1}A_{u} state, sharp autoionizing spectral lines with the line width of Δν =0.13-0.20 cm^{−1} were obtained except that where near the ionization threshold, extensive diffuseness (Δν = 2.1- 4.1 cm^{−1}) was observed. The prominent autoionizing features have been assigned as two Rydberg series converging to the vibronic state (K^{+} =0^{−}, V_{4}^{+} =1) of C_{2}H_{2}^{+}(X̃^{2}Π_{u}). However, if a single rotational level (J′ = 2 − 4, e parity) of V_{4}K_{2} of the Ã state was chosen as the doorway state, no autoionizing Rydberg series was possibly assigned. The interchannel coupling between the vibrationally autoionizing state and the ZEKE state explains the vibronic effect of the ion core on the intensity and spectral line shape of the PFI-ZEKE spectrum of acetylene near the ionization threshold.